ZERO ELECTRON KINETIC ENERGY PULSED FIELD IONIZATION (ZEKE) SPECTROSCOPY OF ZnCH_{3} AND MgCH_{3}

\underline{TIMOTHY A. BARCKHOLTZ}, DAVID E. POWERS, AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

The two-color (1 + 1′) REMPI and ZEKE spectra of the ZnCH_{3} and MgCH_{3} radicals have been obtained. The ZnCH_{3} radicals were generated in a supersonic free jet expansion by ArF excimer photolysis of Zn(CH_{3})_{2} while MgCH_{3} was generated by ablation of a Mg rod in the presence of a methyl source, generally either Sn(CH_{3})_{4} or CH_{3}CN. The ZEKE spectra, taken via various intermediate vibronic levels of the Ã ^{2}E state, are used to reveal the vibrational structure of the ZnCH_{3} and MgCH_{3} cations and to measure the adiabatic ionization energy for each radical. These radicals belong to the C_{3v} point group and are therefore subject to a Jahn-Teller interaction in the Ã state, which allows a broader range of vibrational states to be accessible in the LIF and REMPI experiments, thereby enabling access to a broader range of vibrational states for the ions in the ZEKE experiment. Vibrational frequencies and assignments of the ions will be reported.